OpenStudy (anonymous):

If the measurement of a central angle is 5 pi over 6, find the length of its intercepted arc in a circle with a radius of 15 inches.

OpenStudy (***[isuru]***):

use \[S = r \ \theta\] where S =intercepted arc length r= radius \[\theta = \angle \] can u do that ?

OpenStudy (***[isuru]***):

in your problem \[\theta = \frac{ 5 \pi }{ 6 }\] r= 15 inches plug them in to the equation

OpenStudy (anonymous):

so s would equal 39.2?

OpenStudy (***[isuru]***):

yep!